Circular No. 8538

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 9P/TEMPEL N. Biver, D. Bockelée-Morvan, P. Colom, J. Crovisier, and A. Leca-

cheux, Observatoire de Paris; and G. Paubert, Institut de Radioastronomie Millimétrique (IRAM), report: “Comet 9P/Tempel was observed with the IRAM 30-m radio telescope from May 4.8 to 9.0 UT. The HCN J = 3-2 and J = 1-0 lines were detected with mean integrated intensities of 111 ± 8 and 31 ± 4 mK km/s, respectively. Day-to-day variations in the HCN line intensities are observed and reveal periodic variability in HCN production from 5 to 10 × 10^{24} molecules/s. The periodicity is 1.7 days, consistent with the estimated nucleus rotational period. H_{2}S and CH_{3}OH were also marginally detected; CO, CS and H_{2}CO were searched for and undetected. The production-rate ratios, or their upper limits, of these molecules relative to HCN are 5.5, 21, < 110, < 1.1, and < 17, respectively. All lines are blueshifted, suggesting preferential sunward outgassing.”

Total-magnitude and coma-diameter estimates made with 25×100 binoculars: Apr. 28.97 UT, 10.9, 3′ (J. J. Gonzalez, Asturias, Spain); May 5.54, 10.3, – (D. A. J. Seargent, The Entrance, N.S.W.); 26.90, 10.1, 4′ (M. Lehký, Hradec Králové, Czech Rep.); June 5.97, 10.1, 3′.5 (Gonzalez).

SUPERNOVAE 2005br, 2005bs, 2005cb M. Turatto, S. Benetti, and A. Harutyunyan, Istituto Nazionale di As-

trofisica (INAF), Padova; M. Riello, Cambridge Astronomical Survey Unit, Institute of Astronomy, University of Cambridge; E. Cappellaro, INAF, Napoli; M. T. Botticella, INAF, Teramo; E. Mason, European Southern Observatory (ESO), write that that obtained spectroscopy of three southern supernovae with the ANTU Very Large Telescope (+ Fors2) at ESO on May 25.2 UT under poor weather conditions. Based on the reduced spectra (range 380–920 nm, resolution 1.0 nm), SN 2005br in IC 5084 (IAUC 8516) appears to be a type-Ib SN ∼ 40 days past maximum; the spectrum is very reddened and shows a strong interstellar Na I D absorption with EW ∼ 0.25 nm. SN 2005bs (IAUC 8517) is a normal type-Ia supernova about 30 days past maximum. SN 2005cb in NGC 6753 (IAUC 8530) appears to be a type-Ib/c supernova, similar to SN 1997X (Munari et al. 1998, A.Ap. 333,

159) at about 10 days after maximum.

T. Davis, G. Anderson, and B. Schmidt, Australian National University (ANU), report that a spectrogram (range 340–920 nm) of SN 2005cb (cf. IAUC 8530), obtained on May 19.76 UT with the ANU 2.3-m telescope (+ DBS), reveals it to be a type-Ic supernova ≈ 1 week past maximum light.

2005 June 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green